OpenStudy (anonymous):

Can anyone help me with a math problem PLEASE? The regular math group is too crowded. Here's the problem: -12y - 10 = - 6y + 14 Find y.

OpenStudy (anonymous):

-12y - 10 = - 6y + 14 -6y=24 -6y/-6 = 24/-6 y= -4

OpenStudy (anonymous):

y=-4